OpenStudy (unklerhaukus):

I have a question about methods of solving seconds order differential equations;

OpenStudy (unklerhaukus):

(a)\[\frac{\text d^2y}{\text dx^2}+n^2x=0\] (b)\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]

OpenStudy (unklerhaukus):

My text book suggest using this technique 1 (a)\[\frac{\text d^2y}{\text dx^2}+n^2x=0\]\[\text{let }p=\frac{\text dy}{\text dx}\]\[\frac{\text dp}{\text dx}=\frac{\text d^2y}{\text dx^2}\]\[\frac{\text dp}{\text dx}+n^2x=0\]\[\frac{\text dp}{\text dx}=-n^2x\]\[p=-n^2\int x\text dx\]\[p=-n^2\frac{x^2}{2}+c_1\]\[\frac{\text dy}{\text dx}=-n^2\frac{x^2}{2}+c_1\]\[y=\frac{-n^2}{2}\int x^2+c_1\text dx\]\[y=\frac{-n^2x^3}{6}+c_1x+d_1\]\[y+\frac{n^2x^3}{6}+cx+d=0\] _____________________________ (b)\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]\[\frac{\text d^2y}{\text dx^2}=\sec^2x\]\[\text{let }p=\frac{\text dy}{\text dx}\]\[\frac{\text dp}{\text dx}=\frac{\text d^2y}{\text dx^2}\]\[\frac{\text dp}{\text dx}=\sec^2x\]\[p=\int\sec^2 x\text dx\]\[\frac{\text dy}{\text dx}=\tan(x)+c\]\[y =\int\tan x+c\cdot\text dx\]\[y=\ln \left|\sec x\right|+cx+d\]_____________________________

OpenStudy (unklerhaukus):

but this method seams simpler to me (a)\[\frac{\text d^2y}{\text dx^2}+n^2x=0\]\[\frac{\text d^2y}{\text dx^2}=-n^2x\]\[\iint{\text d^2y}=\iint-n^2x\cdot{\text dx^2}\]\[\int\text dy=\int-n^2\frac{x^2}{2}+c_1\cdot\text dx\]\[y=\frac{-n^2x^3}{6}+c_1x+d_1\]\[y+\frac{n^2x^3}{6}+cx+d=0\] ______________________________ (b)\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]\[\frac{\text d^2y}{\text dx^2}=\sec^2x\]\[\iint\text d^2y=\iint\sec^2x\cdot\text dx^2\]%\cdot\text dx\]\[\int \text dy =\int\tan x+c\cdot\text dx\]\[y=\ln \left|\sec x\right|+cx+d\]______________________________

OpenStudy (unklerhaukus):

{ignore that line starting with %}

OpenStudy (unklerhaukus):

is the second method valid. ?

OpenStudy (zzr0ck3r):

n is some constand?

OpenStudy (zzr0ck3r):

and why not laplace transforms on the second?

OpenStudy (unklerhaukus):

yes n is just a constant term

OpenStudy (anonymous):

yes the second form is valid

OpenStudy (zzr0ck3r):

where does d come from, im confused

OpenStudy (zzr0ck3r):

o i c

OpenStudy (unklerhaukus):

d is the second constant of integration

OpenStudy (unklerhaukus):

or do you mean the infintesimal change \(\text d x\)

OpenStudy (zzr0ck3r):

nope, i see now.

OpenStudy (zzr0ck3r):

I would just use laplace but this all seems cool:)

OpenStudy (anonymous):

i think it will be valid for all equations in form: \[d^2y/dx^2+f(x)=0\]

OpenStudy (zzr0ck3r):

your way on a is quite neet, good job

OpenStudy (unklerhaukus):

i cant seam to find any reference to solving the DE's using the second method,

OpenStudy (zzr0ck3r):

yeah no idea, seems like if it would work we would have seen it. lil scetchy but the "proof" is in the pudding

OpenStudy (anonymous):

how about: \[d^2y/dx^2+f(x)g(y)=0\]

OpenStudy (zzr0ck3r):

thats seperable in the same since so it would prob work?

OpenStudy (anonymous):

like: \[d^2y/dx^2+x^2/y^2=0\]

OpenStudy (unklerhaukus):

\[\frac{\text d^2y}{\text dx^2}+f(x)g(y)=0\]\[\frac{\text d^2y}{\text dx^2}=-f(x)g(y)\]\[\frac{\text d^2y}{g(y)}=-f(x){\text dx^2}\] \[\iint\frac{\text d^2y}{g(y)}=\iint-f(x){\text dx^2}\]

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

damn your fast with that equation editor, i was on line two

OpenStudy (unklerhaukus):

i actually am typing by hand

OpenStudy (zzr0ck3r):

I will show this to my dif eq prob on monday:) or just use it in a test, ha

OpenStudy (zzr0ck3r):

how?

OpenStudy (zzr0ck3r):

o I dont know that code ") and im lazy

OpenStudy (unklerhaukus):

prob*prof ?

OpenStudy (zzr0ck3r):

prof*

OpenStudy (zzr0ck3r):

I dont know how to do the editor well but int(int(dy),dy) would this not be y^2/2+cy+d?

OpenStudy (unklerhaukus):

do i have to use convolution to take the Laplace transform of the second equation?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

actually you could just find the laplace of sec^x

OpenStudy (zzr0ck3r):

err sec^2(x)

OpenStudy (zzr0ck3r):

nm that does not look fun

OpenStudy (unklerhaukus):

i cant remember how to find the laplace transform of that

OpenStudy (zzr0ck3r):

by deffinition int(sec(x)^2*e^(-s*t),t,0,infinity)

OpenStudy (zzr0ck3r):

but thats looks very lame

OpenStudy (zzr0ck3r):

concolution is more about the laplace inverse of a product of laplaces i think

OpenStudy (zzr0ck3r):

convalution*

OpenStudy (unklerhaukus):

ha

OpenStudy (zzr0ck3r):

but what about that double integral thing I mentioned?

OpenStudy (unklerhaukus):

um that is what i first though http://openstudy.com/users/unklerhaukus#/updates/4fe4a333e4b06e92b8730228

OpenStudy (zzr0ck3r):

yeah seems a little odd. but hey why not

OpenStudy (zzr0ck3r):

seems a little to nice a coincidence tou got the right answer both times

OpenStudy (zzr0ck3r):

tou = you

OpenStudy (zzr0ck3r):

ahh

OpenStudy (unklerhaukus):

notice how these are different \[\iint\text d^2y = \iint \text d(\text d y)=\int \text dy=y+c\]this is different to\[\iint\text dx^2 = \iint \text dx\text d x=\int x+c \text dx=\frac{x^2}{2}+cx+d\]

OpenStudy (zzr0ck3r):

yeah I see whawt your saying

OpenStudy (unklerhaukus):

is there any example you can think of where my second method fails

OpenStudy (zzr0ck3r):

nope and I tried like 5 of my questions dealing with integrating factors and seperable equations. But it seems like something must be wrong lol. either way nice work

OpenStudy (foolaroundmath):

How would you integrate \[\int\limits\int\limits \frac{d^{2}y}{g(y)} \] in the example that \[\frac{d^{2}y}{dx^{2}} = f(x)g(y)\]

OpenStudy (unklerhaukus):

\[\iint\frac{\text d^2y}{g(y)}=\iint-f(x){\text dx^2}\]for \(g(y)=1/y^{2}\) \(f(x)=x^2\) \[\iint y^2\text d^2 y =-\iint x^2{\text dx^2}\]

OpenStudy (anonymous):

Your method does seem to work, Rhaukus. As far as I can tell, the substitution is absolutely valid by definition. I cannot say with 100% confidence it works all the time, but I am 95% confident it does.

OpenStudy (foolaroundmath):

so what is \[\int\limits\int\limits y^{2}d^{2}y\] I really dont see how I can evaluate that given the fact that this doesnt make any sense to me

OpenStudy (zzr0ck3r):

read that link he posted to me

OpenStudy (zzr0ck3r):

its seems undefined method but seems to work?

OpenStudy (foolaroundmath):

I read that discussion yesterday. I am still in the dark as to how I would integrate y^2d(d(y)).... doesnt make any sense to me at the moment

OpenStudy (unklerhaukus):

\[\iint y^2\text d \text d y =-\int \left(\int x^2{\text dx}\right)\text dx\] let\[u=\text {d}y \]\[\frac{u^2}{2}=y\]\[\text du=\text {dd}y\] \[\int u^2\text du=-\int \frac{x^3}2+c\cdot \text dx\]\[ \frac{u^3}2 =-\frac{x^4}{4} +cx+d\cdot \text dx\] ?

OpenStudy (unklerhaukus):

should be \(u^3/3\) on the LSH

OpenStudy (unklerhaukus):

i mean \(\frac{u^3}{12}\)

OpenStudy (unklerhaukus):

does that make sense to anyone/

OpenStudy (unklerhaukus):

i dont think the equations i just wrote make any sense anmore

OpenStudy (unklerhaukus):

i think ill just stick to solving these by treating them as y absent second order DEs

OpenStudy (anonymous):

hey....both the methods are same as u r integrating two times in these methods.....and so u get the same ans...

OpenStudy (unklerhaukus):

\[\iint \text d^2y\]still dosent feel right

OpenStudy (anonymous):

wat does it mean?????

OpenStudy (unklerhaukus):

i just can't visualize \(\text d^2y=\text {dd}y\) the infi-infinitesimal .

OpenStudy (anonymous):

∬d2y is always equals to y ,because u are finding the double integral of double derivative... and as integration is the inverse of differentiation.....we get the ans.. as y...

OpenStudy (unklerhaukus):

so is this right then? \[\iint y^2\text d \text d y =\frac{y^4}{12}\]

OpenStudy (foolaroundmath):

I am not convinced in this step: u = dy u^2/2 = y Aren't we supposed to like integrate wrt to something ? for LHS I mean in \[\int\limits u = \int\limits dy\] how can can you write \[\int\limits u = u^{2}/2\] considering that you arent integrating u wrt anything.

OpenStudy (foolaroundmath):

sorry for the late reply ... had gone for lunch

OpenStudy (anonymous):

yes...